OpenStudy (anonymous):

Write the repeating decimal, 0.150150 . . . , as the ratio of two integers in lowest terms.

OpenStudy (anonymous):

these are the choices A) 125/1000 B) 150/1000 C) 150/999 D) 50/999 E) 50/333

OpenStudy (kirbykirby):

\[\begin{align} 0.150150\dots&=\frac{150}{1000}+\frac{150}{1,000,000}+\frac{150}{1,000,000,000}+\ldots\\ &=\frac{150}{1000}\left( 1+\frac{1}{1000}+\frac{1}{1,000,000}+\dots\right)\\ &=\frac{150}{1000}\sum_{i=0}^{\infty}\left(\frac{1}{1000}\right)^i \,\,\,\,\,\,\,\,\text{geometric series }\\ &=\frac{150}{1000}\frac{1}{1-\frac{1}{1000}}\end{align}\]

OpenStudy (kirbykirby):

The rest should be simple simplification. And just be sure your answer is in the lowest terms, as mentioned in the instructions!

OpenStudy (anonymous):

this was really helpful, I wrote down the steps myself to practice how to solve it thank you so much @kirbykirby

OpenStudy (kirbykirby):

:) awesome

OpenStudy (anonymous):

thanks :)

OpenStudy (kirbykirby):

np